The perimeter of a basketball court is 84 meters and the length is 6 meters longer than twice the width. What is the length and width?
Answer by waynest(281)

p e r i m e t e r = 84 ; l e n g t h = 2 w + 6 ; w i d t h = w
2 w + 2 [ 2 w + 6 ] = 84
2 w + 4 w + 12 = 84
6 w + 12 − 12 = 84 − 12
6 w = 72
6 w ∕ 6 = 72 ∕ 6
w = 12
c h e c k :
2 [ 12 ] + 2 [ 2 [ 12 ] + 6 ] = 84
24 + 2 [ 24 + 6 ] = 84
24 + 2 [ 30 ] = 84
24 + 60 = 84
84 = 84